OpenStudy (anonymous):

What values of x and y satisfy x + 2y = 6 and 2x + 3y = 4 simultaneously? x = 0 and y = 3 x = 2 and y = 0 x = -10 and y = 8 x = 8 and y = -1

OpenStudy (mrnood):

what methods do oy uknow for simulataneous equations? x+2y=6 2x+3y= 4 Try multiplying the FIRST equation by 2 Post the new equation here

OpenStudy (anonymous):

would it be 2x+4y=12? sorry if im wrong

OpenStudy (mrnood):

that is correct now the reason we multiplied by 2 is that oyu NOW have 2x in BOTH equations 2x+4y=12 2x +3y =4 SO - if oyu subtract those 2 equaitons you will 'eliminate' x

OpenStudy (anonymous):

so would it be 1y=8?

OpenStudy (mrnood):

yes well done 1y = 8 is th same as y=8 so there is only one possible answer BUT if oyu want to solve for x just put y =8 into the first equation - you can then solve for x

OpenStudy (anonymous):

u may use cramer's method too

OpenStudy (mrnood):

There are seveal methods for simultaneous equaitons in this case elimination was best because y=8 came straight out

OpenStudy (mrnood):

OK - maybe not 'the best' - but very effective. Each method has its merits

OpenStudy (anonymous):

so the answer would be C butb if y=8 how would x=-10? if you dont mind me asking

OpenStudy (anonymous):

@MrNood agreed :)

OpenStudy (mrnood):

After you have solved for one of the variables (in this case y) you can just go back to either of the original equations and use that value so in the FIRST original equaiotn: x + 2(8) = 6 hence x

OpenStudy (anonymous):

oh okay thanks @MrNood

OpenStudy (mrnood):

np - you did the work yourself - I just pointed the way